OpenStudy (-welp-):

Algebra 2 help needed

OpenStudy (-welp-):

What is \[\sqrt[n]{q}^{m}\] in rational exponent form? \[\sqrt[n]{q}^{m}= q ^{\frac{ m }{ n }}?\]

OpenStudy (anonymous):

u are correct

OpenStudy (mathmale):

Yes, you have that right. Now, want to apply it?